OpenStudy (anonymous):

1/x-3+4/x ----------- 4/x-1/x-3..............i need to simplify this

OpenStudy (anonymous):

i will guess it is \[\frac{\frac{1}{x-3}+\frac{4}{x}}{\frac{4}{x-1}-\frac{1}{x-3}}\]

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

pretty damned ugly yes? multiply top and bottom by \[x(x-1)(x-3)\] carefully

OpenStudy (anonymous):

or better yet look at the anwer estudier sent. unfortunately steps aren't given in this answer

OpenStudy (anonymous):

im still confuesd

OpenStudy (anonymous):

Let's just take the top piece, the bottom is very similar. 1/x-3+4/x The lcd of (x-3) and x is x(x-3), right?